# Shuffle-R1: Efficient RL framework for Multimodal Large Language Models via Data-centric Dynamic Shuffle

### v3

@@ 4.3 Advantage-based Batch Shuffle
While PTS could mitigate Advantage Collapsing and improve RL training performance, the Rollout Silencing issue remains unresolved. To overcome this issue, we propose Advantage-based Batch Shuffle (ABS) module that dynamically reshapes training batches to prioritize and reinforce high-value samples. Rather than relying on static data flow, ABS adaptively redistributes trajectories within each training batch, enabling more frequent updates to trajectories with high learning utility. Built on top of PTS, it serves to magnify the gradient exposure of informative samples, reshaping the training data distribution to achieve better data utilization and training efficiency.
Denote a data batch:
$B=\{p_{i}^{g}=(o_{i,1}^{g},\hat{A}_{i,1}^{g},o_{i,2}^{g},\hat{A}_{i,2}^{g},q^{g})\}_{i=1\sim M,g=1\sim G},$ (7)
@@
In our ABS module, we first assign an importance weight to each trajectory pair based on the sum of the absolute advantages:
@@
We then define the sampling probability of each pair $p_{j}\in P_{v}$ as:
$\Phi(p_{j})=\frac{W(p_{j})}{{\sum_{k=1}^{|P_{v}|}}W(p_{k})}.$ (9)
Based on the sampling probability, we perform $S$ sub-sampling from original batch $B$, each sub-sampling has a capacity of $T$ pairs ($2T$ trajectories):
@@
The ABS module optimizes the learning process through Advantage-aware shuffling and Inter-sub-batch resampling. It increases the update frequency of trajectories with higher advantages, maintains diversity while reinforcing high-value samples through repeated exposure. Together, these designs transform each batch into a soft-prioritized structure that better reflects training signal utility.
@@ 5.1 Experimental Setup
We first conduct our experiments on Geometry3K dataset (Lu et al. 2021) ($2.1k$ training samples.) and a subset of MMK12 dataset (Meng et al. 2025) containing the same amount of data, to investigate model performance on limited training resources. Then, we conduct experiments on larger data scale with MM-Eureka dataset (Meng et al. 2025), seeking for better model performance. We randomly select $27k$ samples from MM-Eureka dataset and mix them with Geometry3K dataset, forming a total of $30k$ training corpora. All training samples are in free-form format.
We first perform evaluation on in-domain test set of Geometry3K and MMK12. Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al. 2024b), MathVision (Wang et al. 2024), WeMath (Qiao et al. 2024), MathVista (Lu et al. 2023), HallusionBench (Guan et al. 2024) and ChartQA (Masry et al. 2022). These benchmarks span across math reasoning, visual perception, and chart understanding. We use MathRuler to evaluate questions with free-form ground truths and Gemini-2.0-Flash-001 (Deepmind 2025) to evaluate questions with multi-choice ground truths.
@@
+ Ours 55.89(+17.77) 54.63(+4.81) 69.51(+4.32) 81.64(+1.80)
Table 1: Performance of Shuffle-R1 on Geometry3K dataset compared with GRPO and DAPO.
@@
+ Ours 68.78(+16.65) 55.02(+5.20) 69.87(+4.68) 82.60(+2.76)
Table 2: Performance of Shuffle-R1 on K12 dataset compared with GRPO and DAPO.
@@
We use EasyR1 (Yaowei et al. 2025) as our training codebase. We choose Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct (Bai et al. 2025) as base model to verify our method’s generalizability on model scales. Parameters of vision encoder are kept frozen. We set update batch size to 128 and rollout batch size ($G$) to 512. Rollout temperature is set to 1.0 and learning rate is set to 1e61e-6. All experiments are conducted on $8\times$ 80G GPUs. For PTS, we sample 4 trajectory pairs (8 trajectories) from 16 rollouts for each query, striking a balance between training cost and exploration space. For ABS, we set the sub-sampling batch size ($T$) to 256 pairs (512 query-response trajectories) and the shuffle times ($S$) to 8. For evaluation, we set the temperature to 0.5 and report the average pass@1 accuracy of 8 tests to reduce randomness.
@@ 5.2 Main Results
GPT-4o (Achiam et al. 2024) 50.8 30.4 63.8 68.8 55.0 - -
o1 (OpenAI 2024) 57.0 60.3 73.9 - - - -
Gemini-2.0 pro (Deepmind 2025) 67.3 48.1 71.3 - 49.8 - -
Claude-3.7-Sonnet (Anthropic 2025) 52.0 41.3 66.8 72.6 55.4 - -
@@
InternVL-2.5-8B (Chen et al. 2025b) 39.5 17.0 64.5 - 50.1 79.1 -
InternVL-3-8B (Zhu et al. 2025) - 29.3 71.6 - 49.9 86.6 -
Qwen2.5-VL-3B (Bai et al. 2025) 34.8 21.9 58.4 51.7 59.8 73.1 49.9
Qwen2.5-VL-7B (Bai et al. 2025) 42.6 25.8 67.4 63.5 65.2 79.8 57.4
Cold Start + RL
R1-VL-7B (Zhang et al. 2025) 40.1 24.3 62.3 59.8 60.9 76.1 53.9
Vision-R1-7B (Huang et al. 2025) 46.1 - 70.8 - 57.8 83.1 -
R1-OneVision-7B (Yang et al. 2025) 43.0 24.8 61.2 60.6 66.4 77.8 55.2
OpenVLThinker-7B (Deng et al. 2025) 46.4 24.8 69.7 67.2 59.1 78.4 57.6
VLAA-Thinker-7B (Chen et al. 2025a) 48.9 26.3 69.9 67.7 67.5 80.1 60.1
@@
$M$-Eureka-Qwen-7B (Meng et al. 2025) 49.6 27.4 70.6 67.4 66.7 79.0 60.1
MMR1-Math-7B (Leng et al. 2025) 39.2 31.9 71.5 70.7 69.6 82.0 60.8
ThinkLite-VL-7B (Wang et al. 2025b) 45.2 28.0 72.4 69.3 70.2 82.0 61.2
VL-Rethinker-7B (Wang et al. 2025a) 51.7 29.7 72.0 70.1 69.9 79.0 62.1
NoisyRollout-7B-K12 (Liu et al. 2025a) 50.1 28.0 70.9 70.8 70.1 81.4 62.1
Shuffle-R1-Qwen-3B (Ours) 44.2 26.8 70.4 66.5 69.2 79.9 59.5
Shuffle-R1-Qwen-7B (Ours) 53.9 30.0 77.0 72.3 71.0 84.1 64.7
@@
We compare model performance of our method with GRPO (Shao et al. 2024) and DAPO (Yu et al. 2025). As shown in Tab. 1, after training on the Geometry3K dataset, the 3B model trained with our method achieved 47.88% accuracy on in-domain test set, with a 5.2% improvement compared to model trained with GRPO and 2.7% improvement compared to DAPO. Similarly, for the 7B model, our method led to 3.3% improvement over GRPO and 1.4% over DAPO, achieving 55.89% accuracy. The superiority of our framework is further highlighted in out-of-domain benchmarks. For out-of-domain math reasoning task, the average accuracy is improved by 1.96% on 3B model and 1.5% on 7B model compared to GRPO, and surpasses DAPO as well. Similar performance gains are also observed on HallusionBench and ChartQA, where the data gap is even further from training data, showing strong generalizability of our framework. Tab. 2 demonstrates similar performance results on K12 dataset, where consistent performance gains are achieved on both in-domain and out-of-domain benchmarks, highlighting our framework’s generalizability on different data distribution.
@@
We conduct larger scale experiments on MM-Eureka dataset. As shown in Tab. 3, trained with $30k$ selected data from diverse sources for 150 steps, our 7B model exhibits a substantial accuracy gain over the base model (Qwen2.5-VL-7B). Moreover, it outperforms a series of open-source 7B competitors who also adopt RL training strategies, e.g. MM-Eureka with direct RL and VLAA-Thinker with RL after cold-start SFT. Notably, our model achieves competitive or superior performance on several benchmarks compared to leading close-source models, for instance, Claude-3.7-Sonnet (Anthropic 2025) and GPT-4o (Achiam et al. 2024). Under the same setting, our 3B variant also demonstrates strong performance, even outperforming several 7B models on certain benchmarks. These results highlight the superiority of our proposed approach in boosting the training efficiency in reinforcement learning.
@@
The improvement in model performance mainly stems from better training efficiency. Fig. 2(a) in previous section shows that, under the same update rollout size, model trained under our framework significantly outperforms GRPO, indicating more effective utilization of training data. Advantage distribution analysis in Fig. 4 proves that, PTS effectively mitigates Advantage Collapsing by increasing the proportion of large-magnitude advantages. ABS further optimizes the batch composition, enabling the model to focus on more informative trajectories.
Fig. 5(a) and (b) further probe into training dynamics and demonstrate that our framework consistently achieves higher training and validation accuracy, reaching comparable performance as GRPO but with only half the training steps. Moreover, our framework effectively mitigates the issue of “Rollout Silencing” shown in Fig. 5(c), maintaining a high token utilization rate across all training stages. Fig. 5(d) further illustrates the favorable trade-off between training scale and computational cost of our approach, substantially expanding the RL exploration space with minimal additional overhead.
@@ 5.3 Ablation Study
Figure 5: (a): Training accuracy of GRPO and Shuffle-R1. (b): Validation accuracy of GRPO and Shuffle-R1. (c): Token utilization rate of GRPO and Shuffle-R1. (d): Shuffle-R1 achieves better performance with minimal extra time cost.
We conduct ablation experiments on Qwen2.5-VL-3B-Instruct using Geometry3K, focusing on two objectives: (1) assessing the contribution of each component in our framework, and (2) validating the impact of key designs.
@@
47.88 48.70 63.19 77.04
Table 4: Ablation study on effectiveness of PTS and ABS.
Effectiveness of Algorithm Design.
We assess the effectiveness of proposed PTS and ABS modules by successively introduce them to the baseline. As shown in Tab. 4, in the in-domain Geometry3K test set, PTS demonstrates significant performance gains, improving the model accuracy from 42.64% to 46.21% (+3.57%). When ABS is introduced for co-optimization, the model achieves a further 1.67% performance gain, eventually reaching an accuracy of 47.88%. This improvement trend continues in the out-of-domain evaluation scenario: in the math reasoning task, the full setting (PTS + ABS) achieves an absolute performance improvement of 2.0% compared to GRPO (48.70% vs. 46.74%, with 41.71% before post-training). In the ChartQA dataset, which has more significant distributional differences, the joint application of PTS and ABS still maintains an effective performance gain (77.04% vs. 76.20%, with 73.08% before post-training). Our method also demonstrates improved performance in HallusionBench.
Rationality of Pairwise Trajectory Sampling.
One core mechanism of the PTS lies in the structured contrastive sampling scheme. To verify the effectiveness of this bidirectional optimization mechanism, we designed the following experiments: (1) one-way positive sampling (only select trajectories with highest advantage); (2) one-way negative sampling (only select trajectories with lowest advantage); and (3) unbiased random sampling. All settings maintained a consistent sampling ratio (8 valid trajectories from 16 rollouts) to ensure fairness. We disable ABS since it relies on the pairing result of PTS. As shown in Tab. 5, model trained with PTS receives a consistent performance gain on both in-domain and out-of-domain tasks, while one-way positive and one-way negative sampling show a decline in all tasks, demonstrating the effectiveness and rationality of our design. Unbiased random sampling only receives minor improvement over baseline, far behind the effectiveness of PTS.
@@
+ ABS 47.88 48.80 63.19 77.04
Table 5: Ablation study on rationality of PTS and ABS.
Rationality of Advantage-based Batch Shuffle.
ABS introduces Advantage-aware shuffling and Inter-sub-batch resampling to reshape the training batch. To validate their effectiveness, we designed two contrastive experiments: (1) unbiased shuffle: using uniformly distributed sampling weights to perform shuffle strategy. and (2) static reorder: randomly reorder the training batch without shuffle, maintaining the original data distribution. We enable PTS during training as ABS relies on its pairing result. As shown in Tab. 5, model trained with ABS significantly outperforms the contrastive settings. The unbiased shuffle setting even performs worse compared to the PTS-only setting, demonstrating the rationality of ABS. The static reorder setting has no improvement compared to PTS-only setting, as they have the same data distribution.
@@
We investigate the impact of two key hyperparameters in our framework: (1) the sampling ratio ($\alpha$) in PTS, and (2) the shuffle times ($S$) in ABS. For $\alpha$, we fix $S=8$ and test values of 0.25, 0.5, and 0.75 (i.e., selecting 4, 8, and 12 samples from 16 rollouts). As shown in Fig. 6(a), both $\alpha=0.75$ and $\alpha=0.5$ yield strong performance, while $\alpha=0.25$ lags behind. We attribute this to over-pruning, where filtering out rollouts too aggressively may reduce data diversity. We choose $\alpha=0.5$ to strike a balance between signal quality and computational efficiency. For $S$, we fix $\alpha=0.5$ and vary the shuffle times as 4, 8, and 16. Fig. 6(b) shows that performance improves with increasing $S$, but saturates beyond $S=8$. This suggests that moderate resampling enhances data exposure, but too many shuffles may offer diminishing returns.
Figure 6: Ablation on key hyper parameters. (a): Effect of sampling ratio $\alpha$. (b): Effect of shuffle times $S$.

### v4

@@ 3.4 Advantage-based Batch Shuffle
While PTS mitigates Advantage Collapsing and improve RL training performance, the Rollout Silencing issue remains unresolved. To overcome this issue, we propose Advantage-based Batch Shuffle (ABS) module that dynamically reshapes training batches to prioritize and reinforce high-value samples. Rather than relying on static data flow, ABS adaptively redistributes trajectories within each training batch, enabling more frequent updates to trajectories with high learning utility. Built on top of PTS, it serves to magnify the gradient exposure of informative samples, reshaping the training data distribution to achieve better data utilization and training efficiency.
Denote a data batch provided by PTS:
$B=\{p_{i}^{g}:(o_{i,1}^{g},\hat{A}_{i,1}^{g},o_{i,2}^{g},\hat{A}_{i,2}^{g},q^{g})\}_{i=1\sim M,g=1\sim G},$ (6)
@@
In our ABS module, we first assign an importance weight to each pair $p_{j}\in B$ based on the sum of the absolute advantages:
@@
These weights are then normalized to form a sampling distribution $\Phi$ over the entire batch $B$:
$\Phi(p_{j})=\frac{W(p_{j})}{{\sum_{k=1}^{|B|}}W(p_{k})}.$ (8)
Based on the sampling distribution, we perform $S$ sub-sampling from original batch $B$, each sub-sampling has a capacity of $T$ pairs ($2T$ trajectories):
@@
The ABS module optimizes the learning process through Advantage-aware Shuffling and Sub-batch Resampling. It increases the update frequency of trajectories with higher advantages, maintains diversity while reinforcing high-value samples through repeated exposure. Together, these designs transform each batch into a soft-prioritized structure that better reflects training signal utility.
@@ 4.1 Experimental Setup
We first conduct our experiments on Geometry3K dataset (Lu et al., 2021) ($2.1k$ training samples, ‘Geo3K’ for short) and a subset of MMK12 dataset (Meng et al., 2025) containing the same amount of data (‘K12’ for short), to investigate model performance on limited training resources. To assess the scalability and effectiveness on a larger corpus, we then conduct experiments with MM-Eureka dataset (Meng et al., 2025). Specifically, we construct a $30k$-sample training set by combining the full Geo3K dataset with $27k$ randomly selected samples from MM-Eureka. All training samples are in free-form format.
We first perform evaluation on in-domain test set of Geometry3K and MMK12. Further, as RL is famous for its strong generalizability, we evaluate our model’s performance on the following representative visual reasoning benchmarks: MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024), WeMath (Qiao et al., 2024), MathVista (Lu et al., 2023), HallusionBench (Guan et al., 2024) and ChartQA (Masry et al., 2022). These benchmarks span across math reasoning, visual perception, and chart understanding. We use MathRuler to evaluate questions with free-form ground truths and Gemini-2.0-Flash-001 (Deepmind, 2025) to evaluate questions with multi-choice ground truths. More information in Appendix E.
Table 1: Performance of Shuffle-R1 trained on Geometry3K dataset.
@@
+ GSPO 52.83 54.27 69.48 80.96
[…]
Table 2: Performance of Shuffle-R1 trained on K12 dataset.
@@
+ GSPO 67.44 54.77 69.13 82.04
[…]
@@
We use EasyR1 (Yaowei et al., 2025) as our training codebase. We employ Qwen2.5-VL-3B/7B-Instruct (Bai et al., 2025) as base model to verify our method’s generalizability on model scales. Parameters of vision encoder are kept frozen. We set update batch size to 128 and rollout batch size ($G$) to 512. Rollout temperature is set to 1.0 and learning rate is set to 1e61e-6. All experiments are conducted on $8\times$ H800-80G GPUs. For each query, we generate $2N=16$ rollouts. For PTS, we construct $N=8$ pairs and select the top $M=4$ pairs, corresponding to a retention ratio of $\alpha=0.5$, striking a balance between training cost and exploration space. For ABS, we set $T=256$ pairs (512 query-response trajectories) for each sub-sampling batch. We construct shuffled batch by performing $S=8$ rounds of shuffle. Decoding temperature for evaluation is set to 0.5, and we report average pass@1 accuracy of 8 tests to reduce randomness.
@@ 4.2 Main Results
GPT-4o (Achiam et al., 2024) 50.8 30.4 63.8 68.8 55.0 - -
o1 (OpenAI, 2024) 57.0 60.3 73.9 - - - -
Gemini-2.0 pro (Deepmind, 2025) 67.3 48.1 71.3 - 49.8 - -
Claude-3.7-Sonnet (Anthropic, 2025) 52.0 41.3 66.8 72.6 55.4 - -
@@
InternVL-2.5-8B (Chen et al., 2025b) 39.5 17.0 64.5 - 50.1 79.1 -
InternVL-3-8B (Zhu et al., 2025) - 29.3 71.6 - 49.9 86.6 -
Qwen2.5-VL-3B (Bai et al., 2025) 34.8 21.9 58.4 51.7 59.8 73.1 49.9
Qwen2.5-VL-7B (Bai et al., 2025) 42.6 25.8 67.4 63.5 65.2 79.8 57.4
Cold-Start + RL
R1-VL-7B (Zhang et al., 2025) 40.1 24.3 62.3 59.8 60.9 76.1 53.9
Vision-R1-7B (Huang et al., 2025) 46.1 - 70.8 - 57.8 83.1 -
R1-OneVision-7B (Yang et al., 2025) 43.0 24.8 61.2 60.6 66.4 77.8 55.2
OpenVLThinker-7B (Deng et al., 2025) 46.4 24.8 69.7 67.2 59.1 78.4 57.6
VLAA-Thinker-7B (Chen et al., 2025a) 48.9 26.3 69.9 67.7 67.5 80.1 60.1
@@
$M$-Eureka-Qwen-7B (Meng et al., 2025) 49.6 27.4 70.6 67.4 66.7 79.0 60.1
MMR1-Math-7B (Leng et al., 2025) 39.2 31.9 71.5 70.7 69.6 82.0 60.8
ThinkLite-VL-7B (Wang et al., 2025b) 45.2 28.0 72.4 69.3 70.2 82.0 61.2
VL-Rethinker-7B (Wang et al., 2025a) 51.7 29.7 72.0 70.1 69.9 79.0 62.1
NoisyRollout-7B-K12 (Liu et al., 2025a) 50.1 28.0 70.9 70.8 70.1 81.4 62.1
[…]
@@
We compare our method with GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025). On Geometry3K (Tab. 2), our 3B model reaches 47.88% accuracy, outperforming GRPO by 5.2% and DAPO by 2.7%.; the 7B model achieves 55.89%, with gains of 3.3% and 1.4%, respectively. On out-of-domain math reasoning tasks, our method further improves average accuracy by 1.96% (3B) and 1.5% (7B) over GRPO, also surpassing DAPO. Consistent gains are observed on HallusionBench and ChartQA, where data distribution diverges more from training. Similar trends hold on the K12 experiments (Tab. 2), with improvements in both in-domain and out-of-domain settings, underscoring the robustness of our framework across tasks and distributions. Compared with the latest GSPO (Zheng et al., 2025), which replaces token-level importance sampling with sequence-level importance sampling, our method also demonstrates superior performance on both in-domain and out-of-domain tasks. All the performance results above indicate our framework’s generalizability on different data distributions and model scales, highlighting the effectiveness of dynamic RL training paradigm. More results in Appendix C.
@@
We conduct larger scale experiments on MM-Eureka dataset. As shown in Tab. 3, trained with $30k$ selected data from diverse sources for 150 steps, our 7B model exhibits a substantial accuracy gain over the base model (Qwen2.5-VL-7B). Moreover, it outperforms a series of open-source 7B competitors who also adopt RL training strategies, e.g. MM-Eureka with direct RL and VLAA-Thinker with RL after cold-start SFT. Notably, our model achieves competitive or superior performance on several benchmarks compared to leading close-source models, for instance, Claude-3.7-Sonnet (Anthropic, 2025) and GPT-4o (Achiam et al., 2024). Under the same setting, our 3B variant also demonstrates strong performance, even outperforming several 7B models on certain benchmarks. These results highlight the superiority of our proposed approach in boosting the training efficiency in reinforcement learning.
@@
The improvement in model performance mainly stems from better training efficiency. Advantage distribution analysis in Fig. 5 confirms that, PTS effectively mitigates Advantage Collapsing by increasing the proportion of large-magnitude advantages. ABS further optimizes the batch composition, enabling the model to focus on more informative trajectories.
Fig. 7(a) and (b) further probe into training dynamics and demonstrate that our framework consistently achieves higher training and validation accuracy, reaching comparable performance as GRPO with as little as half the training steps. Moreover, our framework effectively mitigates the issue of “Rollout Silencing” shown in Fig. 7(c), maintaining a high token utilization rate across all training stages. Fig. 7(d) further illustrates the favorable trade-off between training scale and computational cost of our approach, uplifting training performance by a large margin.
Fig. 7 illustrates the actual wall-clock GPU time between GRPO and Shuffle-R1. Under the same update size and total training steps, Shuffle-R1 achieves substantially higher train/val accuracy than GRPO in the early training stage. More importantly, the total GPU time of Shuffle-R1 only increases by 4% $\sim$ 7.7% relative to GRPO. When targeting the same accuracy as GRPO, Shuffle-R1 requires roughly half the number of training steps and approximately 60% of the total wall-clock time.
@@ 4.3 Ablation Study
Figure 6: (a) Training accuracy of Shuffle-R1 on Geo3K. (b) Validation accuracy of Shuffle-R1 on Geo3K. (c) Token utilization rate of Shuffle-R1 on Geo3K. (d) Shuffle-R1 achieves better performance with minimal extra time cost.
Figure 7: Wall-Clock Training Curve of Shuffle-R1 compared with GRPO.
We conduct ablation experiments on Qwen2.5-VL-3B-Instruct using Geometry3K, focusing on two objectives: (1) assessing the contribution of each component in our framework, and (2) validating the impact of key designs. More detailed results are reported in Appendix C.
Component-wise Contribution.
We evaluate the effectiveness of PTS and ABS by incrementally adding them to the baseline. As shown in Tab. 5, On in-domain Geometry3K test set, PTS lifts accuracy from 42.64% to 46.21% (+3.57). Adding ABS yields a further +1.67, reaching 47.88%. Similar improvement trends appear on out-of-domain benchmarks: on the math reasoning set, the full setting (PTS + ABS) attains 48.70% vs. 46.74% with GRPO (41.71% before RL training); on ChartQA, we also observe effective performance gain (77.04% vs. 76.20%, with 73.08% before RL training) despite larger distribution shift. Our method also demonstrates improved performance in HallusionBench. To isolate the contribution of ABS, we conducted an experiment where ABS is applied to trajectory pairs formed by uniform random sampling instead of PTS. This setting (Tab. 5 line 4) also demonstrates notable improvement over the baseline, confirming its effectiveness.
Table 4: Ablation on effectiveness of PTS and ABS.
@@
46.82 47.79 62.77 75.12
[…]
Table 5: Ablation on rationality of PTS and ABS.
@@
[…]
Analysis of Pairwise Trajectory Sampling.
One core mechanism of the PTS lies in the structured contrastive sampling scheme. To validate the effectiveness of our bidirectional, contrastive sampling scheme, we compare PTS against three alternative strategies: (1) One-way Positive Sampling (only select trajectories with highest advantage); (2) One-way Negative Sampling (only select trajectories with lowest advantage); and (3) Unbiased Random Sampling. All settings maintained a consistent sampling ratio (8 valid trajectories from 16 rollouts) to ensure fairness. We disable ABS since it relies on the pairing result of PTS. As shown in Tab. 5, model trained with PTS receives a consistent performance gain on both in-domain and out-of-domain tasks, while both one-way positive and negative sampling result in a performance decline even below the GRPO baseline, demonstrating the effectiveness and rationality of our design. Unbiased random sampling only receives minor improvement over baseline, far behind the effectiveness of PTS.
Figure 8: Ablation on key hyper parameters. (a) Effect of different sampling ratio $\alpha$. (b) Effect of different shuffle times $S$.
Analysis of Advantage-based Batch Shuffle.
ABS introduces Advantage-aware shuffling and Inter-sub-batch resampling to reshape the training batch. To validate their effectiveness, we designed two contrastive experiments: (1) Unbiased Shuffle: using uniformly distributed sampling weights to perform shuffle strategy. and (2) Static Reorder: randomly reorder the training batch without sub-sampling, maintaining the original data distribution. We enable PTS during training as ABS relies on its pairing result. As shown in Tab. 5, model trained with ABS significantly outperforms the contrastive settings. The unbiased shuffle setting even performs worse compared to the PTS-only setting, demonstrating significance of advantage-weighting. The static reorder setting has no improvement compared to PTS-only setting, as they have the same data distribution.
@@
We investigate the impact of two key hyperparameters in our framework: (1) the sampling ratio ($\alpha$) in PTS, and (2) the shuffle times ($S$) in ABS. For $\alpha$, we fix $S=8$ and test values of 0.25, 0.5, and 0.75 (i.e., selecting 4, 8, and 12 samples from 16 rollouts). As shown in Fig. 8(a), both $\alpha=0.75$ and $\alpha=0.5$ yield strong performance, while $\alpha=0.25$ lags behind. We attribute this to over-pruning, where filtering out rollouts too aggressively may reduce data diversity. We choose $\alpha=0.5$ for a balance between signal quality and computational efficiency. For $S$, we fix $\alpha=0.5$ and vary the shuffle times as 4, 8, and 16. Fig. 8(b) shows that performance improves with increasing $S$, but saturates beyond $S=8$. This suggests that moderate resampling enhances data exposure, but too many shuffles may offer diminishing returns. More details in Appendix C.
4.4 Extension Experimental Results
Table 6: Extension experiments on Qwen2.5-VL-32B.
Model MathVerse MathVision MathVista WeMath HallBench ChartQA Avg.
Qwen2.5-32B 57.0 38.2 75.4 72.9 71.3 80.7 65.9
+ GRPO 58.4 39.3 77.3 75.9 70.3 83.8 67.4
[…]
Table 7: Extension experiments on Referring Expression Comprehension task.
Model RefCOCO(testA) RefCOCO(testB) RefCOCO+(testA) RefCOCO+(testB) RefCOCOg(test)
Qwen2.5-3B 86.09 75.64 81.71 66.93 72.39
+GRPO 89.90 81.33 85.94 70.97 81.45
[…]
Scaling to 32B Model
We conduct experiments on Qwen2.5-VL-32B to inveritage the scalability and generalizability of Shuffle-R1 on larger models. We train the 32B variant on the selected $30k$ data for 50 steps (to save training resources cost). Results in Tab. 7 demonstrate that Shuffle-R1 also performs well on 32B scale.
Referring Expression Comprehension
For tasks beyond math/visual reasoning, we conduct an experiment on Referring Expression Comprehension (REC). We train Qwen2.5-VL-3B on 60K data randomly selected from RefCOCO/RefCOCOg/RefCOCO+ using a IoU-based soft reward. Results in Tab. 7 show that Shuffle-R1 can be well adapted to soft-reward tasks and beyond.
Table 8: Extension experiments on LLM.
Setting Math12K AIME24 MATH500 GSM8K GPQA Olymp.
Qwen2.5-Math-1.5B-Base
+ GRPO 67.6 10.0 66.0 74.0 25.7 33.3
[…]
Qwen2.5-Math-7B-Base
+ GRPO 74.6 20.0 76.4 84.8 36.3 39.7
[…]
Language-only Reasoning
While our initial analysis focused on the RL training dynamics of MLLMs, we conduct an early extension experiment to validate the feasibility of Shuffle-R1 on LLMs. We train Qwen2.5-Math-1.5B/7B-Base (Yang et al., 2024) on Open-S1 dataset (Dang and Ngo, 2025) for 150 steps and evaluate model performance on Math12K (Hendrycks et al., 2021), AIME24, MATH500 (Lightman et al., 2023), GSM8K (Cobbe et al., 2021), GPQA-Diamond (Rein et al., 2023) and OlympiadBench (He et al., 2024). As shown in Tab. 8, our framework delivers significant improvements on all the benchmarks compared to GRPO, demonstrating its potential effectiveness on text-only LLMs.